OpenStudy (anonymous):

what is the exact value of cot pi/3

OpenStudy (anonymous):

do you know the value of sin or cos or tan pi/3?

OpenStudy (danjs):

\[\cot(\phi) = \frac{ 1 }{ \tan(\phi) } = \frac{ 1 }{ \frac{ \sin(\phi) }{ \cos(\phi)} } = \frac{ \cos(\phi) }{ \sin(\phi) }\]

OpenStudy (danjs):

cos(pi/3) = 1/2 and sin(pi/3) = sqrt(3) / 2